40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3800     Entry time: Wed Oct 27 21:33:42 2010
Author: Suresh 
Type: Update 
Category: Locking 
Subject: Green from the far end re-obtained 

 

The mirror which was moved during the mode matching of PSL light to the MC (ref elog #3791) has been repositioned.  We once again have the green light from the NPRO on the X (south) arm available on the PSL table. 

This light was supposed to be collimated by the two plano convex lenses (f=200mm and f=50mm  ref to elog #3771) but it was converging.  So I moved the f=50mm lens backwards to make the beam collimated.  I checked the beam collimation by introducing an Al coated mirror infront of th PD and diverting the beam temporarily in a free direction.  I could then check the collinearity and collimation of both the green beams over a meter.  After alignment the mirror was removed and the light is now incident on the PD once again.  We can now proceed to look for green beats.

The power from the PSL NPRO was attenuated for the MC locking work of yesterday.  It has now been increased to the maximum by rotating the Half Wave Plate (HWP).  The power after the PSL is now about 450mW  (500mW - 10% picked off for the doubling).

 

 

 

ELOG V3.1.3-